OpenStudy (anonymous):

pleeeeeeease someone help me! find all solutions of the equation in the interval [0,2pi) sin(x-pi/3)-sin(x+pi/3)=3/2

OpenStudy (mani_jha):

Use the following: \[\sin A+\sin B=2\sin(A+B)/2\cos(A-B)/2\] then solve for sinx. Try it, if any problem, tell me.

OpenStudy (mani_jha):

sinA + sin B =2sin(A+B)/2cos(A-B)/2

OpenStudy (anonymous):

im sorry i dont really understand, can u explain it a little more?

OpenStudy (dumbcow):

i would just use the sum and difference formulas again

OpenStudy (kunal):

i think the answer is (7*pi)/6 or 210 degrees.....

OpenStudy (mani_jha):

It would also be good to use sin(A-B)-sin(A+B)=-2sinBcosA A=x B=pi/3 So sin(x-pi/3)-sin(x+pi/3)=-2sinpi/3cosx=-cosx -cosx=3/2 cosx=-3/2 cos x cant be less than -1. So i guess the question is wrong

OpenStudy (dumbcow):

i agree with Mani Jha question may have a typo or there is no solution

OpenStudy (dumbcow):

wait i double checked, and i made a mistake somewhere kunal is right cos x = -sqrt3/2 x = 5pi/6, 7pi/6

OpenStudy (dumbcow):

@mani jha, sin(pi/3) = sqrt3/2

OpenStudy (mani_jha):

oh yeah, sorry for that mistake